# FiDeLiS: Faithful Reasoning in Large Language Model for Knowledge Graph Question Answering

### v1

@@ 3 Experiments
We aim to answer the following research questions during this section: RQ1: How does FiDeLiS compare with existing baselines in terms of accuracy and efficiency on KGQA tasks? RQ2: How does the design of each individual component contribute to the overall performance of FiDeLiS? RQ3: How robustness of FiDeLiS perform under varying conditions?
3.1 Experiment Settings
Datasets & Metrics. We adopt three benchmark KGQA datasets: WebQuestionSP (WebQSP) [54], Complex WebQuestions (CWQ) [40] and CR-LT-KGQA [11] in this work. We follow previous work [29] to use the same training and testing splits for fair comparison over WebQSP and CWQ. Both WebQSP and CWQ can be reasoned using Freebase KGs [^1] https://github.com/microsoft/FastRDFStore. To address the bias in WebQSP and CWQ, which predominantly feature popular entities and there is a likelihood that their data might have been incorporated into the pre-training corpora of LLMs, we tested our methods on CR-LT-KGQA (further discussed in Section 5). We use the complete dataset from CR-LT-KGQA in our experiments, as it comprises only 200 samples. Each of these can be reasoned based on the Wikidata KG [^2] https://www.wikidata.org/wiki/Wikidata:Main_Page. The statistics of the datasets are given in Table 8 and Table 9.
To streamline the KGs, we utilize a subgraph of Freebase by extracting all triples that fall within the maximum reasoning hops from the question entities in WebQSP and CWQ [29]. Similarly, we construct the corresponding sub-graphs of Wikidata for CR-LT-KGQA as well. We assess the performance of the method by analyzing the F1 and Hits@1 metrics for CWQ and WebQSP, and by evaluating the accuracy for CR-LT-KGQA.
Baselines. (1) RoG [29]: embed structure knowledge graph from KGs into the underlying neural networks during the pretraining and fine-tuning process to generate the reasoning path and explanation. (2) ToG [38]: ask LLM to iteratively explore multiple possible reasoning paths on KGs until the LLM determines that the question can be answered based on the current reasoning paths. Our method FiDeLiS follows this paradigm. (3) NSM [12] utilizes the sequential model to mimic the multi-hop reasoning process. (4) KD-CoT [44] retrieves relevant knowledge from KGs to generate faithful reasoning paths for LLMs. (5) DeCAF [55] combines semantic parsing and LLMs reasoning to jointly generate answers, which also reach salient performance on KGQA tasks. The details of each baseline are further discussed in Section G.1.
Implementation Details. We set the default beam width as $4$ and depth as $4$ without specific annotation. We set the $\alpha$ in Eq 4 as $0.3$ to ensure reproducibility.
Backboned LLMs. We assess our approach on closed- and open-source LLMs. For closed-source LLMs, we choose GPT-4-turbo, GPT-3.5-turbo [^3] We use the recent gpt-4-turbo model that is released in 2024-04-09 from https://platform.openai.com/docs/models/gpt-4-turbo-and-gpt-4 and gpt-3.5-turbo model notated as “gpt-3.5-turbo-0125” from https://platform.openai.com/docs/models/gpt-3-5-turbo. to report and compare the results on all datasets. We use Llama-2-13B [42] and Mistral-7B [19] as our open-source LLMs to conduct cost–performance analysis on different datasets. The experiment results of open-source models can be found in Table 10.
3.2 RQ1: KGQA Performance Comparison
Comparison between baseline methods. Table 1 clearly showcases the comparative effectiveness of FiDeLiS, against different baselines. It shows that FiDeLiS outperforms all baselines integrated with gpt-4-turbo, even for strong baselines being trained, such as DeCAF [55] and RoG [29]. Across all prompting-based methods, gpt-4-turbo achieves higher performance compared to gpt-3.5-turbo, especially on the CWQ dataset. It indicates that gpt-4-turbo has a better understanding and processing capability for complex queries. The CR-LT dataset seems more demanding than WebQSP and CWQ, as implied by the consistently lower accuracy scores. However, FiDeLiS shows consistent enhancements compared to other baselines, demonstrating its capability to handle long-tail entities by referring to the knowledge from KGs and processing more complex queries effectively.
@@ 3.2 RQ1: KGQA Performance Comparison
Finetuning - LLM + KG NSM [12] 74.31 - 53.92 - -
CBR-KBQA [6] - - 67.14 - -
DeCAF [55] 82.1 - 70.42 - -
KD-CoT [44] 73.7 50.2 50.5 - -
RoG [29] 83.15 69.81 61.39 56.17 60.32
Prompting - LLM + KG gpt-3.5-turbo ToG [38] 75.13 72.32 57.59 56.96 62.48
@@
Table 2 provides a detailed view of how removing specific components from FiDeLiS impacts its performance. The results emphasize that some features, such as Path-RAG and beam search in DVBS, are particularly significant, especially for tackling complex reasoning tasks on datasets like CWQ. When Path-RAG and beam-search within DVBS are excluded, we observe substantial declines in performance. This underlines their importance in our model’s ability to handle intricate reasoning effectively. Furthermore, the integration of the reasoning and planning steps within DVBS plays a critical role in maintaining robust performance across varied datasets. One aspect highlighted by our study is the role of deductive verification during the beam search phase in DVBS. This feature enhances the model’s ability to delineate and reinforce decision boundaries, thus safeguarding against reasoning errors.
Table 2: Ablation Studies of FiDeLiS, we test the ablation study using model gpt-3.5-turbo-0125. DVBS refers to “Deductive-verification guided Beam Search’. “$\Delta$” refers to the performance gap between each component and the entire method. The top-3 performance gaps on each dataset are highlighted in green, with darker shades indicating significant performance gap.
- Components WebQSP CWQ CR-LT
@@ 3.4 RQ3: Robustness Analysis
Comparison using different embedding backbones. Table 3 demonstrates the performance of the Path-RAG using various embedding models as backbones. It shows that the integration of the Openai-Embedding Model significantly enhances the performance across all datasets, indicating its superior capability in understanding and processing natural language queries, which leads to more effective retrieval. For instance, the Openai-Embedding Model increases the Hits@1 by 13.04% in the WebQSP dataset and 14.73% in the CWQ dataset compared to the base BM25 [34] model. Furthermore, this improvement is consistent and highlights the feasibility and potential of pairing more powerful language models with structured reasoning frameworks to address complex information retrieval and question answering challenges.
@@
Vanilla Retriever w/ BM25 [34] 58.31 48.39 50.73
w/ SentenceBert [33] 62.74 50.14 51.80
w/ E5 [46]) 68.42 52.84 54.31
@@
Path-RAG w/ BM25 [34] 70.34 56.11 58.77
w/ SentenceBert [33] 73.45 58.41 60.45
w/ E5 [46]) 77.93 62.74 65.23
@@
Comparison using different searching widths and depths. To explore the influence of the search width and depth on FiDeLiS’s performance, we conduct experiments under settings with depths ranging from 1 to 4 and widths from 1 to 4. As shown in Figure 2, FiDeLiS’s performance improves with the search depth and width and when the searching depth is larger than 3, FiDeLiS’s performance start to decrease on both WebQSP and CWQ. This is because only a small part of questions in these two datasets have reasoning depths greater than 3, and also FiDeLiS implements stricter rules for terminating searches using deductive reasoning. However, for searching width, we find that the performance is consistent and implies the FiDeLiS’s performance could potentially be improved with the increment of the exploration breadth. Considering the computational cost (which increases linearly with the depth), we set both the default beam width and depth as 4 without specific annotations in all the experiments.
@@
Figure 2: Performance of FiDeLiS with different beam search widths and depths.
Comparison between Path-RAG and Vanilla Retriever. To verify the validity of the Path-RAG compared to vanilla retrievers. We conduct the experiments as shown in Figure 3. For the vanilla retriever, we concatenate each entity with its corresponding relation to form a reasoning step. We then calculated the cosine similarity between the embeddings of each reasoning step and the query embeddings to select candidate paths.
Figure 3: Comparison between Path-RAG and Vanilla Retriever.
@@
We calculate the coverage ratio (CR) of the retrieved reasoning paths and the ground-truth reasoning paths as follows: $\text{CR}=\left(\frac{N_{\text{retrieved}}\cap N_{\text{ground-truth}}}{N_{\text{ground-truth}}}\right)\times 100\%$, where $N_{\text{retrieved}}$ is the number of retrieved reasoning paths candidates, and $N_{\text{ground-truth}}$ is the total number of ground-truth reasoning paths. We observe that compared to the vanilla retriever, Path-RAG achieves a higher CR value and shows better alignment with the ground-truth paths. It demonstrates that our method can better leverage the connections that are often missed by simpler retrieval models. This feature is critical in subsequent LLM to focus on the information of interest and arrive at a correct and reasonable answer.
@@
Table 4: Average depths of the returned reasoning paths. The average depth of ground-truth reasoning paths of WebQSP is 2.3, for CWQ is 3.2, and for CR-LT is 4.7.
@@
Effectiveness of deductive-verification. To verify the effectiveness of deductive-verification, we calculate the average depths of the grounded reasoning path from ToG and our methods in Table 4. It shows that FiDeLiS consistently shows shorter and closer reasoning depths to ground-truth across all datasets compared to ToG. This implies that our method may offer more precise termination signals and potentially more accurate reasoning paths.
Error Analysis regarding Whole Path Generation.
Figure 4: Validation of whole path generation from RoG [29] over WebQSP.
To verify the faithfulness of our stepwise method, we conduct the error analysis regarding the whole reasoning path generation. Specifically, we conduct an analysis of the validity of the whole reasoning path generation using the baseline methods RoG [29]. The result is shown in Figure 4. The definition of the metric validity ratio (VR) is the ratio of reasoning steps that existed in the knowledge graph to the total number of the reasoning steps in the output reasoning path: $\text{VR}=\left(\frac{N_{\text{valid-steps}}}{N_{\text{all-steps}}}\right)\times 100\%$, where $N_{\text{valid-steps}}$ is the number of reasoning steps that existed in the knowledge graph, and $N_{\text{all-steps}}$ is the number of all the reasoning steps in the reasoning path. We note that 33% of the generated reasoning paths are invalid, 18% exhibit formatting errors, and 15% include relations that do not exist in the knowledge graph. This illustrates the limitations of embedding knowledge graphs into base LLM models, particularly in controlling the generation process.
@@
ToG Iranian rial $\to$ finance.currency.countries_used $\to$ Iran $\to$ location.country.form_of_government $\to$ Islamic republic $\to$ government.form_of_government.countries $\to$ Iran Iranian rial $\to$ finance.currency.countries_used $\to$ Iran $\to$ location.country.form_of_government $\to$ Theocracy $\to$ government.form_of_government.countries $\to$ Iran Iranian rial $\to$ finance.currency.countries_used $\to$ Iran $\to$ location.country.form_of_government $\to$ Unitary state $\to$ government.form_of_government.countries $\to$ Iran The results are Iran.
FiDeLiS Iranian rial $\to$ finance.currency.countries_used $\to$ Iran $\to$ location.country.form_of_government $\to$ Islamic republic Iranian rial $\to$ finance.currency.countries_used $\to$ Iran $\to$ location.country.form_of_government $\to$ Theocracy Iranian rial $\to$ finance.currency.countries_used $\to$ Iran $\to$ location.country.form_of_government $\to$ Unitary state The results are Theocracy, Unitary state, Islamic republic.
@@ 4 Related Work
KG-enhanced LLM. KGs have advantages in dynamic, explicit, and structured knowledge representation and techniques combining LLMs with KGs have been studied [32]. Early studies [29, 55] embed structured knowledge from KGs into the underlying neural networks during the pretraining or fine-tuning process. However, the reasoning steps generated by LMs are observed to be prone to errors, which could be non-existent in the knowledge graph, and can subsequently lead to incorrect reasoning during inference. Also, KG embedded in LLM sacrifices its own nature of explainability in knowledge reasoning and efficiency in knowledge updating [17]. Another solutions are to keep the reasoning model as an agent to explore the external structure knowledge source [38, 9, 20, 44] ToG [38] directly employs LLMs to output scores for candidate selections from the KG. This model operates at the relation-entity level of the KG, aiming to identify relevant triples that aid the LLM in making accurate and responsible final answer predictions.
Knowledge Graph-based Question Answering. Conventional embedding-based methods represent the entities and relations in embedding space and design special model architectures (e.g., Key-Value memory networks, sequential models, and graph neural networks) to reason answers [30, 12, 52]. To integrate LLMs for KGQA, retrieval-augmented methods aim to retrieve the relative facts from the KGs to improve the reasoning performance [25, 22]. Recently, UniKGQA [21] which unifies the graph retrieval and reasoning process into a single model with LLMs, achieves STOA performance. Semantic parsing methods convert the question into a structural query (e.g., SPARQL) by LLMs, which can be executed by a query engine to reason the answers on KGs [39, 24]. However, these methods heavily rely on the quality of generated queries. If the query is not executable, no answers will be generated. DECAF [55] combines semantic parsing and LLMs reasoning to jointly generate answers, which also reach salient performance.

### v2

@@ 3 Experiments
During the experiment section, we aim to answer the following research questions: RQ1: How does FiDeLiS perform compare with existing baselines on KGQA tasks? RQ2: How does the design of each individual component contribute to the overall performance of FiDeLiS? RQ3: How robustness of FiDeLiS perform under varying conditions? RQ4: How about the efficiency of FiDeLiS? All experiment settings are detailed in Appendix D due to page constraints.
@@
Finetuning - LLM + KG NSM (He et al., 2021) 74.31 - 53.92 - -
CBR-KBQA (Das et al., 2021) - - 67.14 - -
DeCAF (Yu et al., 2022) 82.1 - 70.42 - -
KD-CoT (Wang et al., 2023a) 73.7 50.2 50.5 - -
RoG (Luo et al., 2024) 83.15 69.81 61.39 56.17 60.32
Prompting - LLM + KG gpt-3.5-turbo ToG (Sun et al., 2023) 75.13 72.32 57.59 56.96 62.48
@@
3.1 RQ1: KGQA Performance Comparison
Table 1 clearly showcases the comparative effectiveness of FiDeLiS, against different baselines. It shows that FiDeLiS outperforms all baselines using gpt-4-turbo model, even for strong baselines being fine-tuned, such as DeCAF (Yu et al., 2022) and RoG (Luo et al., 2024). Across all prompting-based methods, gpt-4-turbo achieves higher performance compared to gpt-3.5-turbo, especially on the CWQ dataset. It indicates that gpt-4-turbo has a better understanding and processing capability for complex queries. The CR-LT dataset seems more challenging compared with WebQSP and CWQ, as implied by the consistently lower accuracy scores. However, FiDeLiS shows consistent enhancements compared to other baselines, demonstrating its capability to handle long-tail entities by referring to the knowledge from KGs and processing more complex queries effectively.
@@ 3.1 RQ1: KGQA Performance Comparison
Table 2 demonstrates the ablation study of FiDeLiS using the gpt-3.5-turbo-0125 model across the WebQSP, CWQ, and CR-LT datasets. It highlights the critical importance of both the Path-RAG and DVBS components. Removing Path-RAG, whether by employing a vanilla retriever or the think-on-graph (ToG) approach (Sun et al., 2023), results in substantial performance declines, particularly a 6.97% drop in Hits@1 on WebQSP and a 6.01% decrease on CWQ, underscoring the necessity of an effective retrieval mechanism. Similarly, ablating parts of DVBS, especially the beam search component, leads to significant reductions, with an 18.97% decrease in Hits@1 on WebQSP and a 13.34% drop on CWQ, indicating that beam search is vital for maintaining high accuracy. Other DVBS subcomponents, such as the deductive verifier and planning, also contribute notably to performance, albeit to a lesser degree. Overall, the findings demonstrate that the integrated Path-RAG and DVBS frameworks are essential for FiDeLiS’s robust performance, with their removal causing marked decreases in accuracy and Hits@1 scores across all evaluated datasets.
Table 2: Ablation Studies of FiDeLiS using model gpt-3.5-turbo-0125. $\Delta$ refers to the performance gap between each component and the entire method. The three largest performance gaps on each dataset are highlighted in green, with darker shades denoting more significant differences.
Ablation Setting Components WebQSP CWQ CR-LT
@@ 3.3 RQ3: Robustness Analysis
Comparison using different embedding backbones. Table 3 demonstrates the performance of the Path-RAG using various embedding models as backbones. It shows that the integration of the Openai-Embedding Model significantly enhances the performance across all datasets, indicating its superior capability in understanding and processing natural language queries, which leads to more effective retrieval. For instance, the Openai-Embedding Model increases the Hits@1 by 13.04% in the WebQSP dataset and 14.73% in the CWQ dataset compared to the base BM25 (Robertson et al., 2009) model. Furthermore, this improvement is consistent and highlights the feasibility and potential of pairing more powerful language models with structured reasoning frameworks to address complex information retrieval and question answering challenges.
@@
Vanilla Retriever w/ BM25 (Robertson et al., 2009) 58.31 48.39 50.73
w/ SentenceBert (Reimers & Gurevych, 2019) 62.74 50.14 51.80
w/ E5 (Wang et al., 2022a) 68.42 52.84 54.31
@@
Path-RAG w/ BM25 (Robertson et al., 2009) 70.34 56.11 58.77
w/ SentenceBert (Reimers & Gurevych, 2019) 73.45 58.41 60.45
w/ E5 (Wang et al., 2022a) 77.93 62.74 65.23
@@
Comparison using different searching widths and depths. To investigate how the search width and depth affect the performance of FiDeLiS, we have performed experiments with settings ranging from depths of 1 to 4 and widths of 1 to 4. The results, presented in Figure 2, show that performance generally improves as search depth and width increase. However, performance begins to decline when the search depth exceeds 3 for both the WebQSP and CWQ datasets. This decline is attributed to the fact that only a small fraction of questions in these datasets require reasoning at depths greater than 3. Additionally, FiDeLiS utilizes stricter rules for ending the search process, which can preemptively terminate misguided searches. Contrastingly, increasing the search width consistently shows potential for improved performance, suggesting benefits from broader exploration. However, considering the computational costs, which rise linearly with depth, we have chosen to set both the default beam width and depth to 4 for all experiments unless stated otherwise. This default setting aims to balance performance gains with computational efficiency.
(a) RD over CWQ
(b) RD over WebQSP
(c) BW over CWQ
(d) BW over WebQSP
@@
Figure 2: Performance (Hits@1) of FiDeLiS with different beam search widths (BW) and reasoning depths (RD) over WebQSP and CWQ.
@@
Comparison between Path-RAG and Vanilla Retriever. To verify the robustness of the Path-RAG compared to vanilla retrievers. We conduct the experiments as shown in Figure 3(a) and 3(b). We calculate the coverage ratio (CR) of the retrieved reasoning paths and the ground-truth reasoning paths as follows: $\text{CR}=\left(\frac{N_{\text{retrieved}}\cap N_{\text{ground-truth}}}{N_{\text{ground-truth}}}\right)\times 100\%$, where $N_{\text{retrieved}}$ is the number of retrieved reasoning paths candidates, and $N_{\text{ground-truth}}$ is the total number of ground-truth reasoning paths. We set the vanilla retriever as the baseline, specifically, we concatenate each entity with its corresponding relation to form a reasoning step. We then calculated the cosine similarity between the embeddings of each reasoning steps and the query embeddings to select candidate paths. We observe that compared to the vanilla retriever, Path-RAG achieves a higher CR value and shows better alignment with the ground-truth paths. It demonstrates that our method can better leverage the connections that are often missed by simpler retrieval models. This feature is critical in subsequent LLM to focus on the information of interest and arrive at a correct and reasonable answer.
(a) Coverage Ratio over CWQ
(b) Coverage Ratio over WebQSP
(c) Error Analysis over WebQSP
Figure 3: (a)-(b): Empirical study regarding the coverage ratio of the retrieved reasoning paths. (c): Error analysis of the validation of whole paths generated from RoG (Luo et al., 2024) over WebQSP.
Error analysis regarding whole path generation. To verify the faithfulness of our stepwise method, we conduct the error analysis regarding the whole reasoning path generation. Specifically, we conduct an analysis of the validity of the whole reasoning path generation using the baseline methods RoG (Luo et al., 2024). The result is shown in Figure 3(c). The definition of the metric validity ratio (VR) is the ratio of reasoning steps that existed in the knowledge graph to the total number of the reasoning steps in the output reasoning path: $\text{VR}=\left(\frac{N_{\text{valid-steps}}}{N_{\text{all-steps}}}\right)\times 100\%$, where $N_{\text{valid-steps}}$ is the number of reasoning steps that existed in the knowledge graph, and $N_{\text{all-steps}}$ is the number of all the reasoning steps in the reasoning path. It show that only 67% of generated reasoning steps are valid, while the remaining 33% of reasoning steps either have a format error or do not exist in the KG. This illustrates that the reasoning steps generated by language models offer few guarantees about feasibility especially when multiple consecutive steps are combined into a reasoning path.
@@
Table 4: Average depths of the returned reasoning paths. GT refers to ground-truth reasoning paths.
@@
Effectiveness of deductive-verification. To verify the effectiveness of deductive-verification, we calculate the average depths of the grounded reasoning path from ToG and our methods in Table 4. It shows that FiDeLiS consistently shows shorter and closer reasoning depths to ground-truth across all datasets compared to ToG. This implies that our method may offer more precise termination signals and potentially more accurate reasoning paths compared to baselines.
@@
ToG Path-1: Iranian rial $\to$ finance.currency.countries_used $\to$ Iran $\to$ location.country.form_of_government $\to$ Islamic republic $\to$ government.form_of_government.countries $\to$ Iran Path-2: Iranian rial $\to$ finance.currency.countries_used $\to$ Iran $\to$ location.country.form_of_government $\to$ Theocracy $\to$ government.form_of_government.countries $\to$ Iran Path-3: Iranian rial $\to$ finance.currency.countries_used $\to$ Iran $\to$ location.country.form_of_government $\to$ Unitary state $\to$ government.form_of_government.countries $\to$ Iran Based on the reasoning paths, the result is Iran.
FiDeLiS Path-1: Iranian rial $\to$ finance.currency.countries_used $\to$ Iran $\to$ location.country.form_of_government $\to$ Islamic republic Path-2: Iranian rial $\to$ finance.currency.countries_used $\to$ Iran $\to$ location.country.form_of_government $\to$ Theocracy Path-3: Iranian rial $\to$ finance.currency.countries_used $\to$ Iran $\to$ location.country.form_of_government $\to$ Unitary state Based on the reasoning paths, the results are Theocracy, Unitary state, Islamic republic.
3.4 RQ4: Efficiency of FiDeLiS
To investigate the runtime efficiency and cost efficiency of FiDeLiS, we present a comparison regarding the average runtime, average token usage, average times of LLM calling per question in Table 6. We found that (1) our method shows superior efficiency compared to the ToG (which is also training-free), by reducing approximately 1.7x runtime costs. (2) Path-Rag component is critical in enhancing both the accuracy and efficiency of the model. Its ability to constrain potential path candidates effectively reduces unnecessary computational overhead, leading to quicker and more accurate results.
Table 6: Runtime efficiency of FiDeLiS per question.
Dataset Method Hits@1 (%) Avg Runtime (s) Avg Token Usage Avg LLM calling
WebQSP FiDeLiS (ours) 79.32 43.83 2,452 10.7
   w/o Path-RAG using vanilla retriever 72.35 48.37 2,873 10.7
   w/o Path-RAG using ToG 75.11 74.26 6,437 10.7
FiDeLiS (ours) - GPT-4o 81.17 37.82 2,452 10.7
FiDeLiS (ours) - GPT-4o-mini 76.48 24.31 2,452 10.7
CWQ FiDeLiS (ours) 63.12 74.59 2,741 15.2
   w/o Path-RAG using vanilla retriever 57.11 78.41 3,093 15.2
   w/o Path-RAG using ToG 59.47 132.59 5,372 15.2
FiDeLiS (ours) - GPT-4o 65.33 50.12 2,741 15.2
FiDeLiS (ours) - GPT-4o-mini 58.34 42.54 2,741 15.2
To address concern regarding our method’s potential application in real-time scenarios, we also tested our method using faster and more advanced LLMs. Table 6 shows that our method could be further accelerated with newer, faster models like GPT-4o or GPT-4-mini. The potential of the ongoing advancements in LLMs are expected to further enhance the scalability and efficiency of FiDeLiS, making it a practical development in challenging environments. More detailed analysis of bottleneck of computation of FiDeLiS can be further found in Appendix C.
@@ 4 Related Work
KG-enhanced LLM. KGs have advantages in dynamic, explicit, and structured knowledge representation and techniques combining LLMs with KGs have been studied (Pan et al., 2024). Early studies (Luo et al., 2024; Yu et al., 2022) embed structured knowledge from KGs into the underlying neural networks during the pretraining or fine-tuning process. However, the reasoning steps generated by LMs are observed to be prone to errors, which could be non-existent in the knowledge graph, and can subsequently lead to incorrect reasoning during inference. Also, KG embedded in LLM sacrifices its own nature of explainability in knowledge reasoning and efficiency in knowledge updating (Hu et al., 2023). Another solutions are to keep the reasoning model as an agent to explore the external structure knowledge source (Sun et al., 2023; Gu et al., 2023; Jiang et al., 2023b; Wang et al., 2023a) ToG (Sun et al., 2023) directly employs LLMs to output scores for candidate selections from the KG. This model operates at the relation-entity level of the KG, aiming to identify relevant triples that aid the LLM in making accurate and responsible final answer predictions.
Knowledge Graph-based Question Answering. To integrate LLMs for KGQA, retrieval-augmented methods aim to retrieve the relative facts from the KGs to improve the reasoning performance (Li et al., 2023; Karpukhin et al., 2020). Recently, UniKGQA (Jiang et al., 2022) which unifies the graph retrieval and reasoning process into a single model with LLMs, achieves STOA performance. Semantic parsing methods convert the question into a structural query (e.g., SPARQL) by LLMs, which can be executed by a query engine to reason the answers on KGs (Sun et al., 2020; Lan & Jiang, 2020). However, these methods heavily rely on the quality of generated queries. If the query is not executable, no answers will be generated. DECAF (Yu et al., 2022) combines semantic parsing and LLMs reasoning to jointly generate answers, which also reach salient performance.